OpenStudy (anonymous):

The graph of g(x) is f(x) translated to the left 8 units and up 2 units. What is the function rule for g(x) given f(x)=x^2 Please show steps. Thank you! Will give a medal to best answer.

OpenStudy (jdoe0001):

f(x) = A ( Bx + C ) + D C = horizontal shift, C > 0, to the left, C < 0, to the right D = vertical shift, D > 0, up, D < 0, down

OpenStudy (jdoe0001):

so, what do you think our C and D will be?

OpenStudy (anonymous):

I'm not sure at all how to figure these types of problems out.

OpenStudy (jdoe0001):

notice f(x) = A ( Bx + C ) + D C = horizontal shift, C > 0, to the left, C < 0, to the right D = vertical shift, D > 0, up, D < 0, down our C and D components are all we need :)

OpenStudy (jdoe0001):

---> translated to the left 8 units and up 2 units <----

OpenStudy (anonymous):

C>8 D>2

OpenStudy (jdoe0001):

if it's translated/shifted, to the left, that is horizontal shift, that means C >0, or positive so +8 if it's shifted up, by 2 units, that means D > 0, or positive, that means +2 our "parent function" is \(\bf x^2\) so, to shift it \(\bf f(x) = x^2\\\quad \\ \textit{shifted to the left by }8 \qquad \implies (x+8)^2\\\quad \\ \textit{pelleted up by }2 \qquad \implies (x+8)^2+2\)

OpenStudy (jdoe0001):

pelleted? heheh,

OpenStudy (jdoe0001):

lol I have a typo.... hhehe

OpenStudy (jdoe0001):

\(\bf f(x) = x^2\\\quad \\ \textit{shifted to the left by }8 \qquad \implies (x+8)^2\\\quad \\ \textit{shifted up by }2 \qquad \qquad \implies (x+8)^2+2\\\quad \\ g(x) = (x+8)^2+2\)

OpenStudy (anonymous):

Ooh lol thank you